OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (foreverandalways13):

hang on!

OpenStudy (anonymous):

https://www.youtube.com/watch?v=6FmWAvyTYBA&feature=endscreen&NR=1 this may help

OpenStudy (ivettef365):

you have x - 4y = 1, which is the same as x = 4y +1 now substitute on the other equation the x for 4y+1 and find value of y

OpenStudy (foreverandalways13):

I got this guys! lol

OpenStudy (foreverandalways13):

or use elimination...its so much easier..lol

OpenStudy (foreverandalways13):

y=-1 plug that in to an original equation:) do I need to explain how I got y?

OpenStudy (anonymous):

(3, 1)

OpenStudy (anonymous):

?

OpenStudy (foreverandalways13):

I think so let me check

OpenStudy (foreverandalways13):

-3

OpenStudy (foreverandalways13):

but yeah u got this!

OpenStudy (anonymous):

6.01] Select the system of equations that corresponds to the given graph. A coordinate plane shows two lines on the graph. One line crosses the y axis at (0, 2) and passes through the point (2,1). The other line crosses the y axis at (0,-2) and passes through the point (2,-1). -2x - y = 4 2x + y = -2 -8x + 4y = -16 -4x + 2y = -8 x + 2y = 4 2x - 4y = 8 2x + 3y = 6 3x + 2y = 4 --------------------------------------------------------------------------------

OpenStudy (foreverandalways13):

bleh..this might take a while:) but in reality u really could plug in those points for each equation...

OpenStudy (foreverandalways13):

it takes time but it works:) make sure it works for BOTH equations...

OpenStudy (foreverandalways13):

I could do this if I wasn't doing psychology too:P I have 4 days to finish like 7 units...BLEH!!!

OpenStudy (foreverandalways13):

well that helps! I think ur answer is where the lines cross...

OpenStudy (foreverandalways13):

haha a little late on the graph thing! lol

OpenStudy (anonymous):

computer was being stuped

OpenStudy (foreverandalways13):

oh yeah that happens...lol

OpenStudy (anonymous):

x + 2y = 4 2x - 4y = 8

OpenStudy (foreverandalways13):

ummm hang on! I got this

OpenStudy (foreverandalways13):

its infinitely many...